OpenStudy (anonymous):

Find the accumulated value of an investment of $5000.00 at 5% compounded monthly for 8 years.

OpenStudy (wolf1728):

Total = Principal * (1 + (rate /n))^n*years where n = 12 for 12 months per year Total = 5,000 * (1 + (.05/12))^12*8 Total = 5,000 * (1.0041666667)^96 Total = 5,000 * 1.4905854727 Total = 7,452.93 And here's a calculator to check your answer

OpenStudy (anonymous):

thank you!!!

OpenStudy (wolf1728):

Total = Principal * (1 + (rate /n))^n*years where n = 12 for 12 months per year Total = 5,000 * (1 + (.05/12))^12*8 Total = 5,000 * (1.0041666667)^96 Total = 5,000 * 1.4905854727 Total = 7,452.93 And here's a calculator to check your answer http://www.1728.org/compint.htm You are welcome :-)